OpenStudy (anonymous):

The coordinate of the midpoint of segment AB is (-3, 2). Find the coordinate of A if the coordinate of B is (0, 2). A = (-6,2) (-3/2,2) (3,2)

OpenStudy (anonymous):

@alyisnumber1 @nikato

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@babygirl19961216

OpenStudy (anonymous):

@Bad.Coding @Juicstice

OpenStudy (e.mccormick):

Know how the midpoint formula is done?

OpenStudy (anonymous):

is it (-3/2,2)?

OpenStudy (e.mccormick):

For the x or y value of a midpoint, it is the average of the other points. So the average of the xes is the midpoint x. So you would plug in the known and solve for the unknown. I have not done the process, so not sure here. There is another way as well. A little logic about what the midpoint means and how much change is happening.

OpenStudy (e.mccormick):

For any line from \((x_1,y_1)\) to \((x_2,y_2)\) the midpoint \((x_m,y_m)\) is \(\left ( \dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2} \right )\). So put in what you know and solve for the unknown with: \(x_m=\dfrac{x_1+x_2}{2}\) \(y_m=\dfrac{y_1+y_2}{2}\)

OpenStudy (anonymous):

so its (3,2)? im confused

OpenStudy (e.mccormick):

Well, what are you getting as the setup when you plug values into one of those?